OpenStudy (anonymous):

2x^2 Does the ^2 apply only to the x or 2 and x

OpenStudy (anonymous):

just to the "x"

OpenStudy (anonymous):

for example \[2\times 3^2=2\times 9=18\] not \[2\times 3^2=6^2=36\]

OpenStudy (anonymous):

so if x= 4, then how would this work

OpenStudy (anonymous):



OpenStudy (anonymous):

it would be \[2\times 4^2=2\times 16=32\]

OpenStudy (anonymous):

32

OpenStudy (anonymous):

satellite beat me to it!

OpenStudy (anonymous):

Alright Thanks you Guys!